A certain recipe ask for 3/8 cup of granola. How much is needed to make 7/9 of the recipe?
Answer by stanbon(75887)

A certain recipe ask for 3/8 cup of granola. How much is needed to make 7/9 of the recipe?
-----------------------
Ans: (7/9)(3/8) = (7/3)(1/8) = 7/24 cup
-------------------------
Cheers,
Stan H.
--------------------